# AI-Enhanced 7-Point Checklist for Melanoma Detection Using Clinical Knowledge Graphs and Data-Driven Quantification

### v1

@@ Abstract
The 7-point checklist (7PCL) is widely used in dermoscopy to identify malignant melanoma lesions needing urgent medical attention. It assigns point values to seven attributes: major attributes are worth two points each, and minor ones are worth one point each. A total score of three or higher prompts further evaluation, often including a biopsy. However, a significant limitation of current methods is the uniform weighting of attributes, which leads to imprecision and neglects their interconnections. Previous deep learning studies have treated the prediction of each attribute with the same importance as predicting melanoma, which fails to recognize the clinical significance of the attributes for melanoma. To address these limitations, we introduce a novel diagnostic method that integrates two innovative elements: a Clinical Knowledge-Based Topological Graph (CKTG) and a Gradient Diagnostic Strategy with Data-Driven Weighting Standards (GD-DDW). The CKTG integrates 7PCL attributes with diagnostic information, revealing both internal and external associations. By employing adaptive receptive domains and weighted edges, we establish connections among melanoma’s relevant features. Concurrently, GD-DDW emulates dermatologists’ diagnostic processes, who first observe the visual characteristics associated with melanoma and then make predictions. Our model uses two imaging modalities for the same lesion, ensuring comprehensive feature acquisition. Our method shows outstanding performance in predicting malignant melanoma and its features, achieving an average AUC value of 85%. This was validated on the EDRA dataset, the largest publicly available dataset for the 7-point checklist algorithm. Specifically, the integrated weighting system can provide clinicians with valuable data-driven benchmarks for their evaluations.
{IEEEkeywords}
Melanoma, 7-point checklist, data-driven, classification, multi-modal
1 Introduction
[…]
@@ 1 Introduction
Skin cancer is a widespread malignancy worldwide, with melanoma being the most dangerous type. Although the incidence of skin cancer varies by region and individual risk factors, timely diagnosis and prompt intervention can significantly improve the prognosis of the disease [me2024global].
@@
In the medical field, healthcare professionals utilize diverse imaging modalities to expedite the diagnosis and treatment of skin lesions [heibel2020review]. Two notable techniques are dermoscopic imaging and clinical photography. Dermoscopy, which is an optical magnification technique assisted by liquid immersion and cross-polarized illumination, produces intricate images revealing subepidermal structures and lesions to support the diagnosis [celebi2019dermoscopy]. On the other hand, clinical images efficiently capture essential morphological features including color, texture, shape, and borders, despite their inability to explore intricate subcutaneous lesion details. The convenience and accessibility of clinical images, especially through widely available mobile electronic devices like smartphones, are indisputable benefits [wen2022characteristics]. The fusion of these two imaging techniques provides a comprehensive information essential in practical medical scenarios, specifically for melanoma diagnosis.
Figure 1: Example of biopsy-confirmed melanoma cases from the EDRA dataset, accompanied by attributes outlined in the 7-point checklist.
@@
Conventionally, clinical practice relies on pattern analysis, which recognizes observed dermatologic attributes to aid in skin lesion diagnosis. Several established algorithms, such as the ABCD rule [abbasi2004early] and the 7-point checklist (7PCL) [kawahara2018seven], aid dermatologists in recognizing cancer-related characteristics, contributing to a simpler diagnostic process. For example, this study focuses on the 7PCL algorithm, which evaluates seven dermoscopic characteristics associated with malignant melanoma. These include three major attributes: atypical pigment network (ATP-PN), blue-whitish veil (BWV), and irregular vascular structures (IR-VS), as well as four minor attributes: irregular pigmentation (IR-PIG), irregular streaks (IR-STR), irregular dots and globus (IR-DaG), and regression structures (RS). Each major attribute receives a weighted score of two points, while minor attributes earn one point each. Cumulative scores establish an initial evaluation and necessitate a predetermined threshold, generally established at one or three points, as shown in Fig. 1. While the 7PCL algorithm has gained widespread acceptance and usage, limitations exist in the traditional method due to subjectivity resulting from the user’s experience and lack of clarity regarding the respective weights of major and minor factors. Subjective adjustments to thresholds can also impact the sensitivity and specificity of the algorithm. Therefore, implementing objective quantitative criteria can enhance predictive performance, benefiting healthcare practitioners and facilitating the training of professionals in the field.
1.1 Related Work
@@
Computer-aided diagnosis (CAD) has rapidly emerged as an effective tool to assist dermatologists in the detection of melanoma, as it can extract characteristic information related to the lesion. The conventional CAD procedure for melanoma diagnosis based on machine learning involves multiple steps. These include image preprocessing (e.g., artifact removal and color correction), lesion segmentation, extraction of hand-crafted feature sets, and classification using dedicated classifiers based on the application context [barata2017development, celebi2007methodological, schaefer2014ensemble, fabbrocini2014automatic, wadhawan2011implementation]. Initially, Celebi et al. [celebi2007methodological] introduced a method that automatically detects borders, extracts shape, color, and texture features from relevant regions, selects optimized features, and addresses class imbalance specifically in dermoscopy images of pigmented skin lesions. Based on this work, researchers have proposed ways to take advantage of the variety of hand-crafted features of melanoma. Schaefer et al. [schaefer2014ensemble] proposed an ensemble method that combines multiple classifiers and various hand-crafted features and demonstrated statistically improved recognition performance. Fabbrocini et al. [fabbrocini2014automatic] proposed a framework that integrates hand-crafted features and the 7-point checklist. The authors employed both machine learning classifiers and statistical analysis to enhance the accuracy of diagnosis. Wadhawan et al. [wadhawan2011implementation] devised a comparable notion and modified it for intelligent, portable devices aimed at routine clinical skin monitoring or as an aid for dermatologists and primary care physicians in identifying melanoma. These efforts have facilitated collaboration between machine learning methods and clinical diagnosis of melanoma. However, their performance has been limited by an over-reliance on preprocessing and complicated feature extraction procedures.
@@
Over the past decade, deep learning has become a widely used and powerful tool for feature learning and pattern recognition in CAD, enabling automated detection of melanoma and other skin diseases in various scenarios using different methods [esteva2017dermatologist, wang2021deep, abbas2019dermodeep, barata2018survey, abhishek2021predicting, barata2023reinforcement]. Esteva et al. [esteva2017dermatologist] demonstrated the classification of skin lesions using a single convolutional neural network trained end-to-end directly from both clinical and dermoscopy images, using only pixels and disease labels as input, and achieving performance comparable to dermatologists. Then, Abhishek et al. [abhishek2021predicting] presented the first work to directly explore image-based predicting clinical management decisions of melanoma without explicitly predicting diagnosis. Meanwhile, the International Skin Imaging Collaboration (ISIC) challenges and the corresponding constantly updated and expanded public datasets have greatly contributed to the rapid development of the field, and more excellent methods and training strategies have been proposed [codella2018skin]. However, the above-mentioned research works focus more on the pixel information brought by the image itself, especially in the imaging modality of dermoscopy. Incorporating expert domain knowledge and multimodal data are important directions that deserve further attention.
@@
Using information from clinical diagnostics to support multi-modal deep learning models has proven effective in several related research. Moura et al. [moura2019abcd] combined the ABCD rules and a pre-trained CNNs through a multi-layer perceptron classifier for melanoma detection and achieved improved accuracy. Kawahara et al. [kawahara2018seven] proposed a multitask Convolutional Neural Network (CNN) trained on both dermoscopy and clinical images, along with the corresponding 7PCL attributes for melanoma detection, which has been regarded as the benchmark of the related research studies. Following this work, Bi et al. [bi2020multi] proposed a hyper-connected CNN (HcCNN) structure for the multi-modal analysis of melanoma and the corresponding clinical features. Tang et al. [tang2022fusionm4net] further developed this idea and fused 7PCL attributes information into a deep learning framework in a multi-stage manner, improving the average diagnostic accuracy. However, none of these studies have fully considered the intrinsic connections between melanoma attributes when analyzing the clinical data. The meta-information is processed consistently as feature vectors in combination with image features, resulting in some improvement in the outcomes. Nevertheless, the potential value of this clinical information warrants further investigation. To address these issues, Graph Convolutional Neural Networks (GCNs) are introduced to assist with data mining for works involving supplementary clinical information. Wu et al. [wu2020learning] introduced GCNs based on general clinical meta-information to aid in the multi-label classification of skin diseases. Wang et al. [wang2021incorporating] designed a framework utilizing constrained classifier chains, and first examined the mutual relationships between attributes of 7PCL. Fu et al. [fu2022graph] applied GCNs to a 7PCL-based multimodal and multilabeled classification task, and uncovered the interconnections between attributes by analyzing simultaneous occurrences of different attributes and undirected GCNs. These studies demonstrate the potential of combining graph learning with clinical diagnosis to aid dermatologists in the pre-diagnosis of melanoma. However, a gap persists in understanding the directional relationships among diverse attributes and their practical implications for melanoma diagnosis. This is crucial for fully realizing the benefits of rule-based clinical algorithms like 7PCL.
1.2 Contributions
@@
[…]
  • We propose a graph module for the 7PCL clinical algorithm, focusing on disease diagnosis results as the central node and attributes as peripheral nodes. This module separates multi-labeled information into internal and external interactions, enabling joint learning through trainable hyperparameters. Additionally, we introduce directed weight relationships between labeled nodes, transforming clinical experience into data-driven visual relationships to enhance predictive ability and practicality.
  • The method explores not only the typical first- and second-order connections of graph networks but also incorporates connection information from higher-order sensory domains. This encompasses indirect relationships between individual nodes along extended paths during graph network construction. Consequently, it addresses the imbalance between isolated nodes and receptive domains, yielding more comprehensive information for multi-label feature extraction.
  • We develop an end-to-end melanoma diagnostic system based on the 7PCL algorithm. Our system predicts melanoma by assigning trainable weights to each attribute, followed by a decision layer using the 7PCL attributes. This framework maintains clinical diagnosis logic while introducing new data-driven criteria, aiding dermatologists, general practitioners, and trainees in both practice and education.
  • We evaluate our method on the Interactive Atlas of Dermoscopy (EDRA), showing improved prediction accuracy by enhancing deep learning models’ use of meta-information. For generalization, we also test on ISIC 2017 and ISIC 2018 datasets, confirming similar attribute-based improvements.
Figure 3: Illustration of our proposed melanoma diagnosis framework, which incorporates a topological graph convolutional neural network inspired by clinical knowledge and a gradient diagnosis mechanism with a data-driven weighted module that mimics the behavior of dermatologists. In the framework, (a) section represents a multimodal fusion framework that combines comprehensive information from two imaging modalities used in this work. Here, D refers to dermoscopy images, C refers to clinical images, and F refers to the fused feature after the max pooling function. The (b) section shows the generation of the graph information learned from the attributes listed in the 7PCL. Within this section, both internal and external mutual connections between melanoma and its related attributes are considered. The generated 7PCL graph feature embeddings are fed to the deep features from imaging modalities. Section (c) introduces a gradient diagnostic strategy that elaborates on the joint optimization functions for both attributes and melanoma. The trainable weights module serves as the bridge for this gradient workflow.
2 Methodology
@@
As illustrated in Fig. 3, the proposed method comprises three modules. The first module, Clinical-Dermoscopic Multimodal Fusion (CD-MFM), combines information from clinical and dermoscopic modalities. The second module, 7-Point Checklist Directed Graph Mining (7PCL-DirGM), extracts representative features using a mining strategy that leverages directed and multi-order graph information from multi-labeled data. These two modules form our Clinical Knowledge-Based Topological Graph Convolutional Network (CKTG). Additionally, the GD-DDW employs data-driven weighted gradient diagnosis to enhance diagnostic accuracy by emulating a dermatologist’s diagnostic behavior. The following subsections provide a detailed overview of the methodology.
2.1 Clinical-Dermoscopic Multimodal Fusion (CD-MFM)
@@
In this study, we treated each skin lesion as an individual “case” within the dataset. Each case, denoted as ${x}^{i}$, encompasses data from multiple modalities, including dermoscopy images ${x}^{i}_{\textrm{d}}$, clinical images $x^{i}_{\textrm{c}}$, and the encoded information related to the $j_{\textrm{th}}$ attribute within the 7PCL, $x^{i}_{\textrm{7p}_{j}}$, where $i$ ranges from 1 to $n$ (the total number of cases) and $j$ ranges from 1 to 7. Additionally, each case is associated with a diagnostic tag, denoted as $y^{i}_{\textrm{m}}$, which represents the melanoma diagnosis, and labels $y^{i}_{\textrm{7p}_{j}}$ for their morphologic attributes.
To commence, a deep convolutional neural network was applied to extract visual features independently from dermoscopic and clinical images. Utilizing ResNet-50 as the backbone model yielded feature vectors $X_{\textrm{c}}$ and $X_{\textrm{d}}$. Following this, multimodal feature fusion was conducted by combining feature channels extracted from dermoscopic and clinical images through weighted averaging, resulting in an additional channel denoted as $X_{\textrm{f}}$.
$X_{\textrm{f}}={\delta X_{\textrm{c}}+(1-\delta)X_{\textrm{d}}}$ (1)
@@
Let $Z$ represent the features extracted from the graph information. For the next fusion step, we individually combined the graph features ($Z$) with the deep features extracted from dermoscopic ($X_{\textrm{d}}$), clinical ($X_{\textrm{c}}$), and the additional channel ($X_{\textrm{f}}$) images. This was achieved by element-wise multiplication:
$X_{\textrm{fused\_d}}$ $=X_{\textrm{d}}\odot Z$ (2)
$X_{\textrm{fused\_c}}$ $=X_{\textrm{c}}\odot Z$
$X_{\textrm{fused\_f}}$ $=X_{\textrm{f}}\odot Z$
@@
where $X_{\textrm{fused\_d}}$, $X_{\textrm{fused\_c}}$, and $X_{\textrm{fused\_f}}$ represent the fused features for the dermoscopic, clinical, and additional channel images, respectively.
@@
Subsequently, we apply a fully connected layer to each channel independently and conduct a weighted average fusion of the features:
$X=\gamma_{\textrm{d}}X_{\textrm{fused\_d}}+\gamma_{\textrm{c}}X_{\textrm{fused\_c}}+\gamma_{\textrm{f}}X_{\textrm{fused\_f}}$ (3)
@@
The resulting fused feature $X$ with dimensions $\mathbb{R}^{n\times d}$ was used for downstream tasks, where $n$ denotes the number of cases and $d$ represents the feature dimensionality, leveraging the combined information from the 7PCL graph and the image modalities.

### v3

# Integrating Clinical Knowledge Graphs and Gradient-Based Neural Systems for Enhanced Melanoma Diagnosis via the 7-Point Checklist

@@ Abstract
The 7-point checklist (7PCL) is a widely used diagnostic tool in dermoscopy for identifying malignant melanoma by assigning point values to seven specific attributes. However, the traditional 7PCL is limited to distinguishing between malignant melanoma and melanocytic Nevi, and falls short in scenarios where multiple skin diseases with appearances similar to melanoma coexist. To address this limitation, we propose a novel diagnostic framework that integrates a clinical knowledge-based topological graph (CKTG) with a gradient diagnostic strategy featuring a data-driven weighting system (GD-DDW). The CKTG captures both the internal and external relationships among the 7PCL attributes, while the GD-DDW emulates dermatologists’ diagnostic processes, prioritizing visual observation before making predictions. Additionally, we introduce a multimodal feature extraction approach leveraging a dual-attention mechanism to enhance feature extraction through cross-modal interaction and unimodal collaboration. This method incorporates meta-information to uncover interactions between clinical data and image features, ensuring more accurate and robust predictions. Our approach, evaluated on the EDRA dataset, achieved an average AUC of 88.6%, demonstrating superior performance in melanoma detection and feature prediction. This integrated system provides data-driven benchmarks for clinicians, significantly enhancing the precision of melanoma diagnosis.
Index Terms:
graph convolutional neural networks, data-driven learning, multi-modal, multi-label, skin cancer detection
I Introduction
@@
Skin cancer is a widespread malignancy worldwide, with melanoma being the most dangerous type. Although the incidence of skin cancer varies by region and individual risk factors, timely diagnosis and prompt intervention can significantly improve the prognosis of the disease [1].
@@
In the medical field, healthcare professionals utilize diverse imaging modalities to expedite the diagnosis and treatment of skin lesions [2]. Two notable techniques are dermoscopic imaging and clinical photography. Dermoscopy, which is an optical magnification technique assisted by liquid immersion or cross-polarized illumination, produces intricate images revealing subepidermal structures and lesions to support the diagnosis [3]. On the other hand, clinical images efficiently capture essential morphological features including color, texture, shape, and borders, despite their inability to explore intricate subcutaneous lesion details. The convenience and accessibility of clinical images, especially through widely available mobile electronic devices like smartphones, are indisputable benefits [4]. The fusion of these two imaging techniques provides a comprehensive information essential in practical medical scenarios, specifically for melanoma diagnosis.
@@
Conventionally, clinical practice relies on pattern analysis, which recognizes dermatologic attributes to aid in skin lesion diagnosis. Several established algorithms, such as the ABCD rule [5] and the 7-point checklist (7PCL) [6], aid dermatologists in recognizing cancer-related characteristics, contributing to a simpler diagnostic process. For example, this study focuses on the 7PCL algorithm, which evaluates seven dermoscopic characteristics associated with malignant melanoma. These include three major attributes: atypical pigment network (ATP-PN), blue-whitish veil (BWV), and irregular vascular structures (IR-VS), as well as four minor attributes: irregular pigmentation (IR-PIG), irregular streaks (IR-STR), irregular dots and globus (IR-DaG), and regression structures (RS). Each major attribute receives a weighted score of two points, while minor attributes earn one point each. Cumulative scores provide an initial evaluation, requiring a predetermined threshold—commonly set at one or three points. Scores exceeding this threshold are classified as suspected melanoma, as illustrated in Fig. 1. Although the 7PCL algorithm is widely accepted and utilized, the attributes and their assigned weights are primarily designed to differentiate between malignant melanoma (Mel) and melanocytic Nevi (MN) [7], limiting its range of applications. By analyzing these attributes as key features using more appropriate quantitative criteria, there is potential to broaden its applicability and enhance predictive performance, ultimately benefiting healthcare practitioners.
In this study, we propose a data-driven graph-based framework for melanoma diagnosis that incorporates clinical insights from the 7-point checklist (7PCL). Our work is based on the observation that melanoma attributes exhibit directed relationships, influencing each other to varying degrees, as shown in Fig. 2. Additionally, the relative importance of these attributes varies depending on the diagnostic context. In summary, the main contributions of this work are:
  • We develop a directed weighted GCN encoding clinically relevant relationships between melanoma attributes, transforming expert knowledge into a structured, data-driven diagnostic system.
  • A dual-attention mechanism enables an effective fusion of image features and meta-information, improving feature extraction and predictive performance.
  • Our method assigns trainable weights to 7PCL attributes, ensuring adaptability across different dermatological conditions while preserving the clinical decision logic.
  • Our method is evaluated in three open access datasets, EDRA, ISIC2017, ISIC2018, demonstrating improved melanoma prediction and attribute recognition.
Figure 1: Example of biopsy-confirmed melanoma cases from the EDRA dataset, annotated with attributes from the 7-point checklist: (a) IR-PIG (1), IR-STR (1), IR-DaG (1), BWV (2), and ATP-PN (2), totaling 7 points; (b) IR-DaG (1), IR-VS (2), BWV (2), and IR-PIG (1), totaling 6 points; (c) IR-DaG (1), ATP-PN (2), IR-PIG (1), and RS (1), totaling 5 points. All cases exceed the threshold of 3 points.
II Related Work
@@
Computer-aided diagnosis (CAD) has rapidly emerged as an effective tool to assist dermatologists in the detection of melanoma, as it can extract characteristic information related to the lesion. The conventional CAD procedure for melanoma diagnosis based on machine learning involves multiple steps. These include image preprocessing (e.g., artifact removal and color correction), lesion segmentation, extraction of hand-crafted feature sets, and classification using dedicated classifiers based on the application context [8, 9]. Initially, Celebi et al. [10] introduced a method that automatically detects borders, extracts shape, color, and texture features from relevant regions, selects optimized features, and addresses class imbalance specifically in dermoscopy images of pigmented skin lesions. Based on this work, researchers have proposed ways to take advantage of the variety of hand-crafted features of melanoma. Schaefer et al. [11] proposed an ensemble method that combines multiple classifiers and various hand-crafted features and demonstrated statistically improved recognition performance. Fabbrocini et al. [12] proposed a framework that integrates hand-crafted features and the 7-point checklist. The authors employed both machine learning classifiers and statistical analysis to enhance the performance of diagnosis. Wadhawan et al. [9] devised a comparable notion and modified it for intelligent, portable devices aimed at routine clinical skin monitoring or as an aid for dermatologists and primary care physicians in identifying melanoma. These efforts have facilitated collaboration between machine learning methods and clinical diagnosis of melanoma. However, their performance has been limited by an over-reliance on preprocessing and complicated feature extraction procedures.
@@
Over the past decade, deep learning has become a widely used and powerful tool for feature learning and pattern recognition in CAD, enabling automated detection of melanoma and other skin diseases in various scenarios using different methods [13, 14, 15, 16]. Esteva et al. [13] demonstrated the classification of skin lesions using a single convolutional neural network trained end-to-end directly from both clinical and dermoscopy images, using only pixels and disease labels as input, and achieving performance comparable to dermatologists. Then, Abhishek et al. [15] presented the first work to directly explore image-based predicting clinical management decisions of melanoma without explicitly predicting diagnosis. Meanwhile, the International Skin Imaging Collaboration (ISIC) challenges and the corresponding constantly updated and expanded public datasets have greatly contributed to the rapid development of the field, and more excellent methods and training strategies have been proposed [17]. However, the above-mentioned research works focus more on the pixel information brought by the image itself, especially in the imaging modality of dermoscopy. Incorporating expert domain knowledge and multimodal data are important directions that deserve further attention.
@@
Using information from clinical diagnostics to support multi-modal deep learning models has proven effective in several related research. Moura et al. [18] combined the ABCD rules and a pre-trained CNNs through a multi-layer perceptron classifier for melanoma detection and achieved improved performance. Kawahara et al. [6] proposed a multitask Convolutional Neural Network (CNN) trained on both dermoscopy and clinical images, along with the corresponding 7PCL attributes for melanoma detection, which has been regarded as the benchmark of the related research studies. Following this work, Bi et al. [19] proposed a hyper-connected CNN (HcCNN) structure for the multi-modal analysis of melanoma and the corresponding clinical features. Tang et al. [20] further developed this idea and fused 7PCL attributes information into a deep learning framework in a multi-stage manner, improving the average diagnostic performance. However, none of these studies have fully considered the intrinsic connections between melanoma attributes recorded as meta-information when analyzing the clinical data. The meta-information is processed consistently as feature vectors in combination with image features, resulting in some improvement in the outcomes. Nevertheless, the potential value of this clinical information warrants further investigation. To address these issues, Graph Convolutional Neural Networks (GCNs) are introduced to assist with data mining for works involving supplementary clinical information [21, 22, 23]. Wu et al. [24] introduced GCNs based on general clinical meta-information to aid in the multi-label classification of skin diseases. Wang et al. [25] designed a framework utilizing constrained classifier chains, and first examined the mutual relationships between attributes of 7PCL. Fu et al. [22] applied GCNs to a 7PCL-based multimodal and multilabeled classification task, and uncovered the interconnections between attributes by analyzing simultaneous occurrences of different attributes and undirected GCNs. These studies demonstrate the potential of combining graph learning with clinical diagnosis to aid dermatologists in the pre-diagnosis of melanoma. However, a gap persists in understanding the directional relationships among diverse attributes and their practical implications for melanoma diagnosis.
In recent studies, deep learning algorithms incorporating attention mechanisms have gained prominence and achieved significant success in medical image and signal analysis [26, 27, 28, 29]. In the field of dermatology, He et al. [30] introduced a cross-modality structure that applies attention weight maps learned separately from bimodal images to each other, enhancing prediction performance. Xu et al. [31] developed an architecture that integrates global and local attention mechanisms, mimicking the diagnostic process of dermatologists. Additionally, Xiao et al. [32] proposed an attention-guided feature reconstruction method utilizing both clinical and ultrasound images, leveraging intra- and inter-modality information to extract more discriminative features. While these studies have significantly enhanced the performance of deep learning models for dermatological image analysis and improved the interpretability of their results, they primarily focus on interactions among learned image features alone. The integration of meta-information, such as the 7PCL attributes, and its interaction with image features remains underexplored. Furthermore, the potential for AI technology to provide feedback on the 7PCL and thereby benefit clinical decision-making has yet to be addressed.
@@
Figure 3: Illustration of our proposed melanoma diagnosis framework, featuring a topological graph convolutional neural network inspired by clinical knowledge and a gradient diagnosis mechanism. The framework includes: (a) the multimodal fusion system combining data from dermoscopy and clinical images, and the meta information; (b) generation of graph features from the 7PCL attributes; and (c) the gradient diagnostic strategy with a trainable weights module.
III Methodology
@@
As illustrated in Fig. 3, the proposed method comprises three modules. The first module, Clinical-Dermoscopic Multimodal Fusion (CD-MFM), combines information from clinical and dermoscopic modalities. The second module, 7-Point Checklist Directed Graph Mining (7PCL-DirGM), extracts representative features using a mining strategy that leverages directed and multi-order graph information from multi-labeled data. These two modules form our Clinical Knowledge-Based Topological Graph Convolutional Network (CKTG). Additionally, the GD-DDW employs data-driven weighted gradient diagnosis to enhance diagnostic performance by emulating a dermatologist’s diagnostic behavior. The following subsections provide a detailed overview of the methodology.
Figure 4: Illustration of the dual-attention blocks
III-A Clinical-Dermoscopic Multimodal Fusion (CD-MFM)
@@
In this study, we treated each skin lesion as an individual “case” within the dataset. Each case, denoted as ${x}^{i}$, encompasses data from multiple modalities, including dermoscopy images ${x}^{i}_{\textrm{d}}$, clinical images $x^{i}_{\textrm{c}}$, and the encoded information related to the $j_{\textrm{th}}$ attribute within the 7PCL, $x^{i}_{\textrm{7p}_{j}}$, where $i$ ranges from 1 to $n$ (the total number of cases) and $j$ ranges from 1 to 7. Additionally, each case is associated with a diagnostic tag, denoted as $y^{i}_{\textrm{m}}$, which represents the melanoma diagnosis, and labels $y^{i}_{\textrm{7p}_{j}}$ for their 7PCL attributes.
First, residual convolutional blocks are applied to independently extract visual features from dermoscopic and clinical images, denoted as $X_{\textrm{d}}$ and $X_{\textrm{c}}$, respectively. Simultaneously, a BERT-based semantic embedding layer is used to encode the label information into semantic features $X_{\textrm{m}}$, which represent the diagnosis and 7PCL attributes. These semantic features are aligned with the visual features from the two image modalities. Next, two separate concatenation layers are applied to combine the visual features and the semantic features, resulting in $X_{\textrm{dm}}$ for the dermoscopic modality and $X_{\textrm{cm}}$ for the clinical modality. These combined features are then passed through dual-attention feature extraction blocks for further processing.
As shown in Fig. 4, the dual-attention feature extraction blocks consist of both Single-Modality Attention (SMA) [33] and Cross-Modality Attention (CMA) blocks [30]. The features $X_{\textrm{dm}}$ and $X_{\textrm{cm}}$ are first refined by their respective SMA blocks, which capture intra-modality interactions within the dermoscopic and clinical image embeddings, embedding_D and embedding_C, respectively, along with their meta-information in embedding_m. These features then pass through residual blocks before being processed by the CMA block, which facilitates cross-modality interaction between the dermoscopic and clinical features. This enables the integration of complementary information from both modalities, enhancing the joint feature representation.
In detail, as shown in Fig. 5, the input features $X_{\textrm{dm}}$ and $X_{\textrm{cm}}$ are transformed into query ($Q$), key ($K$), and value ($V$) representations for each modality. The attention scores for each modality are then computed as:
$S_{\textrm{dm}}$ $=\text{softmax}\left(\frac{Q_{\textrm{dm}}\cdot K_{\textrm{dm}}^{T}}{\sqrt{d_{k}}}\right)$ (1)
$S_{\textrm{cm}}$ $=\text{softmax}\left(\frac{Q_{\textrm{cm}}\cdot K_{\textrm{cm}}^{T}}{\sqrt{d_{k}}}\right)$
The refined features for each modality are then obtained by applying the attention scores to the corresponding value representations in the SMA block:
$X_{\textrm{SMA\_dm}}$ $=S_{\textrm{dm}}\cdot V_{\textrm{dm}}$ (2)
$X_{\textrm{SMA\_cm}}$ $=S_{\textrm{cm}}\cdot V_{\textrm{cm}}$
Similarly, the CMA block facilitates interaction between the two modalities by sharing attention across them. The cross-modality refined features are computed as:
$X_{\textrm{CMA\_dm}}$ $=S_{\textrm{dm}}\cdot V_{\textrm{cm}}$ (3)
$X_{\textrm{CMA\_cm}}$ $=S_{\textrm{cm}}\cdot V_{\textrm{dm}}$
where $X_{\textrm{CMA\_dm}}$ represents the dermoscopic features refined using clinical information, and $X_{\textrm{CMA\_cm}}$ represents the clinical features refined using dermoscopic information. By leveraging both intra- and cross-modality attention, the model ensures that relevant information from both dermoscopic and clinical images is fully captured and integrated for a more robust melanoma diagnosis.
Next, as shown in Fig. 3, after the feature extraction process through the four dual-attention blocks, we obtain the refined feature embeddings $X_{\textrm{d\_refined}}$ and $X_{\textrm{c\_refined}}$ for the individual image modalities, which are combined with meta-information extracted via the SMA module. Additionally, the fusion feature $X_{\textrm{f\_refined}}$ is obtained through the CMA module. These three components are then prepared for further fusion with the information provided by the graph neural network 7PCL-DirGM.
@@
Let $Z$ represent the features extracted from the graph information. In the next fusion step, we individually combine the graph features ($Z$) with the deep features using element-wise multiplication:
$X_{\textrm{d\_fused}}$ $=X_{\textrm{d\_refined}}\odot Z$ (4)
$X_{\textrm{c\_fused}}$ $=X_{\textrm{c\_refined}}\odot Z$
$X_{\textrm{f\_fused}}$ $=X_{\textrm{f\_refined}}\odot Z$
@@
where $X_{\textrm{d\_fused}}$, $X_{\textrm{c\_fused}}$, and $X_{\textrm{f\_fused}}$ represent the fused features for the dermoscopic, clinical, and combined channels, respectively.
(a) SMA Block
(b) CMA Block
Figure 5: The illustration of Single-Modality Attention (SMA) block and Cross-Modality Attention (CMA) block. The operator $\oplus$ denotes element-wise addition, while $\otimes$ represents element-wise multiplication.
@@
Subsequently, we apply a fully connected layer to each channel independently and perform a weighted average fusion of the features:
$X=\gamma_{\textrm{d}}X_{\textrm{d\_fused}}+\gamma_{\textrm{c}}X_{\textrm{c\_fused}}+\gamma_{\textrm{f}}X_{\textrm{f\_fused}}$ (5)
@@
The resulting fused feature, $X$, with dimensions $\mathbb{R}^{n\times 768}$, is utilized for downstream tasks, where $n$ represents the number of cases and 768 corresponds to the feature dimensionality, aligned with the meta-embedding layer. This process effectively integrates information from both the 7PCL graph and the image modalities.
1